OpenStudy (anonymous):

solve by substitution method 0.1x+0.9y=-4.1 0.6x-0.7y=5.9

OpenStudy (anonymous):

this time is your turn. Try to find the answer by yourself. The first thing to to here is to convert 0.1, 0.9, 4.1, 0.6, 0.7 and 5.9 into a fraction. 0.1 = 1/10 0.9=9/10 4.1=41/10 0.6=6/10 0.7=7/10 5.9=59/10 You have to follow the procedure Mathmind gave you before. If you have any question, ask :)

OpenStudy (anonymous):

im stuck on this one with the decimals. I tried the formula but something is not adding up right.

OpenStudy (anonymous):

the equation after the conversion is: (1/10)x+(9/10)y=-41/10 -> equation1 (6/10)x-(7/10)y=59/10 -> equation2 now choose one equation (either 1 or 2) and isolate one variable (4 example x). Try to isolate x from equation 1

OpenStudy (anonymous):

1/10x+9/10y=41/10

OpenStudy (anonymous):

Because both equation 1 and 2 have the same denominator, you can simplify it. The equation system will look like this: x+9y=-41 -> equation1 6x-7y=59 -> equation2 Now isolating x from equation one, we get: x=-41-9y What is the next step to take?

OpenStudy (anonymous):

I need your feedback man :)

OpenStudy (anonymous):

thinking

OpenStudy (anonymous):

6(-41-9y)-7y=59

OpenStudy (anonymous):

Exactly! If we isolated x from equation 1, we have to replace this x (-41-9y) into equation 2. This is like a rule. For example, if you isolate x from equation 2, then you have to replace this x into equation 1. Now find the value for y

OpenStudy (anonymous):

-246-45y-7y=59 -246-52y=59 -52y=305

OpenStudy (anonymous):

oops math error i multiplied wrong

OpenStudy (anonymous):

ok redo it and show me the correct answer.

OpenStudy (anonymous):

y=5

OpenStudy (anonymous):

exactly you got it. Now you have to replace this value (y=5) into either eq1 or 2 and you will find the x value. That is all you have to do my friend. Well done :) Sorry for my English :( my native language is spanish

OpenStudy (anonymous):

im inserting the 5 for y but its not adding up right

OpenStudy (anonymous):

let me check it out

OpenStudy (anonymous):

y is not equal to 5, but -5

OpenStudy (anonymous):

try with this value for y

OpenStudy (anonymous):

Do you understand what substituting means?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

I know you know but Im asking rich if he understands the concept of substituting.

OpenStudy (anonymous):

no but im trying to

OpenStudy (anonymous):

Well from the question you know that you are trying to solve for x and y but both of them have x and y in it which can be completely different so you want to essentially get a single variable x or y it doesnt matter in one equation and you solve for that one.

OpenStudy (anonymous):

that is exactly what we have been doing :)

OpenStudy (anonymous):

How can he understand what you are doing when he doesn't understand the concept. If you don't understand something and you work it out and find the answer it does no help since he is just memorizing it.

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Yeah you are right romero. We had a discution about that before :). Rich you finally did it. If you replace x and y into the one of the original equation you have both sides are equal. Well done dude

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

That is the procedure you have to follow

OpenStudy (anonymous):

np